### v1

@@ Abstract
The rise of autonomous vehicles (AVs) promises to significantly enhance transportation safety and efficiency by mitigating human error, which is responsible for over 90% of road accidents. However, the increasing connectivity of AVs introduces new cybersecurity challenges, as traditional perimeter-based security models are inadequate for dynamic and untrusted environments. This paper presents a novel Zero Trust-based Decentralized Identity Management (D-IM) protocol for AVs. By integrating the core principles of Zero Trust Architecture, "never trust, always verify", with the tamper resistant and decentralized nature of a blockchain network, our framework eliminates reliance on centralized authorities and provides continuous verification for every entity. We detail the system’s design, which leverages Hyperledger Iroha to enable lightweight and secure authentication without a central trusted entity. A comprehensive experimental evaluation, conducted across both urban and highway scenarios, validates the protocol’s practicality. Our results demonstrate that the D-IM framework introduces minimal overhead, with less than 7.5% reduction in Packet Reception Rate (PRR) in urban settings and an increase of under 11% in Channel Busy Ratio (CBR) for LTE-V2X. These findings prove the protocol’s efficiency and robustness, providing a resilient foundation for securing real-time V2X communication against impersonation and replay attacks.
{IEEEkeywords}
@@ 1 INTRODUCTION
[…]
Autonomous Vehicles (AVs) are revolutionizing modern transportation by offering improved road safety, efficiency, and convenience, while minimizing human error that is responsible for more than 90% of road accidents and more than 40,000 fatalities and 2 million injuries annually in the United States alone (Figure 1) [1, 2]. AVs integrate sophisticated sensors, Electronic Control Units (ECUs), and advanced communication protocols to enable semi- and fully autonomous driving. These systems rely on Global Positioning Systems (GPS), LiDAR, cameras, and Vehicle-to-Everything (V2X) communication for perception, situational awareness, and decision-making [3].
The Society of Automotive Engineers (SAE) defines six levels of driving automation, ranging from Level 0 (no automation) to Level 5 (full automation). These levels describe the degree of vehicle autonomy in terms of steering control, environment monitoring, and fallback responsibility [4, 5]. Table 1 summarizes the SAE autonomy levels. As AVs progress toward higher autonomy (aiming for Level 5), their reliance on connectivity becomes indispensable. V2X communication, including Vehicle-to-Vehicle (V2V), Vehicle-to-Infrastructure (V2I), and Vehicle-to-Network (V2N) enables cooperative mobility. Key use cases include platooning, where vehicles form convoys to reduce drag and improve fuel efficiency; drone swarms, coordinating with AVs for delivery or traffic monitoring; emergency vehicle prioritization, where AVs yield dynamically to ambulances and police; and cooperative perception, where vehicles share sensor data for beyond-line-of-sight awareness [6]. Dedicated Short-Range Communications (DSRC) and cellular-V2X (C-V2X) are the primary enabling radio access technologies, supporting low-latency and high-reliability communication.
@@ 5 Full S S S All
Figure 1: Different Reasons for Road Accidents.
In platooning, for instance, the lead vehicle periodically broadcasts kinematic states (e.g., position, speed, acceleration) and intent, while the follower vehicles fuse these messages with onboard sensors to compute control inputs that maintain inter-vehicle gaps and string stability where the disturbances are not amplified when propagating along the vehicles string. This increasing interconnection introduces severe cybersecurity concerns. Malicious actors can impersonate a leader vehicle, injecting false acceleration or braking commands that endanger the convoy. Identity management thus becomes critical for validating trust in dynamic vehicular environments. Current solutions primarily rely on Public Key Infrastructure (PKI) and pseudonym certificates, but these approaches face significant challenges in credential revocation, often failing to achieve timely and efficient invalidation of compromised credentials, thereby exposing vehicles to prolonged risk [7]. Recent reviews highlight that secure and scalable identity management remains one of the most pressing V2X research challenges [6, 5].
The increased connectivity and complexity of AVs increases AVs cybersecurity vulnerabilities introducing many attack vectors. Attack vectors such as GPS spoofing, sensor manipulation, V2X spoofing and information injection threaten vehicle safety and operational integrity. Real-world incidents, such as the Jeep Cherokee and Tesla Model S exploits, have demonstrated the potential for remote hijacking and malicious manipulation of vehicle systems [8]. These incidents highlight the inadequacy of traditional perimeter-based security models, which assume trust within a protected network, and the urgent need for robust multilayered defense strategies [9]. Unfortunately, many traditional approaches rely on trusted entities for credential management, a critical weakness adversaries can exploit by stealing or maliciously issuing credentials.
@@
Consider a platoon of autonomous vehicles traveling in tight formation on a highway, continuously exchanging messages to coordinate acceleration, braking, and lane changes. While this cooperative driving strategy improves safety and fuel efficiency, it also introduces significant vulnerabilities. A malicious vehicle may attempt to infiltrate the platoon by impersonating a legitimate member. Once inside, the adversary could launch attacks such as fake data injection, transmitting falsified distance or braking information that destabilizes the convoy, or GPS spoofing, misleading the platoon about its true location and causing unsafe maneuvers. These manipulations not only threaten the cohesion of the platoon but also heighten the risk of chain-reaction accidents. Such scenarios highlight the necessity of robust identity verification and trust management mechanisms to safeguard cooperative vehicular systems.
Zero Trust Architecture (ZTA) addresses limitations of perimeter models by eliminating implicit trust and enforcing continuous verification of all entities, whether inside or outside the network. ZTA aligns with AV needs where real-time decisions depend on authenticated, accurate data, and is advocated by NIST for modern, interconnected systems [10].
Within this paradigm, identity management emerges as the core security enabler for V2X. Ensuring that only authenticated, authorized, and non-revoked entities can publish or act upon safety messages is essential, especially in dense, dynamic traffic. Existing solutions, notably Public key infrastructure (PKI) with short-term pseudonym certificates, face well-documented shortcomings: reliance on centralized CAs, privacy–linkability trade-offs, scalability under high churn, and slow or bandwidth-heavy revocation distribution that jeopardizes timeliness [11, 7].
Existing Identity management approaches suffer from centralization, scalability bottlenecks, and slow or bandwidth-heavy revocation mechanisms where a single trusted authority with implicit trust is needed presention a potential single point of failure [11, 7]. To address these problems, we design a Zero Trust–based Decentralized Identity Management (D-IM) system. Our solution eliminates dependence on centralized authorities, supports (through the use of blockchain technology) low-latency identity proofing that meets V2X timing budgets, enables fast and tamper-evident revocation propagation across dynamic topologies, and scales to dense traffic scenarios while preserving privacy through unlinkable pseudonym use. By embedding identity verification into a Zero Trust framework, the design ensures that every entity in the network is continuously authenticated and authorized before contributing safety-critical data [5, 10].
@@
This paper presents a Zero Trust–based Decentralized Identity Management (D-IM) system for AVs, instantiated on a permissioned blockchain (Hyperledger Iroha). D-IM leverages blockchain-backed identifiers and verifiable credentials to bind cryptographic identities to vehicles and infrastructure, enforces continuous verification at enrollment and message time, and supports fast, tamper-evident revocation. Conceptually, every entity, vehicle, sensor, or infrastructure node, must prove authenticity before joining the network. Identity proofs accompany safety messages so consumers can validate origin, freshness, and authorization before actuation. Voting-based cross-validation further reduces single-source risk in cooperative applications. The decentralized ledger provides immutable auditability and eliminates implicit trust in perimeter boundaries, while Zero Trust policies restrict privileges to the minimum necessary and re-evaluate them continuously. Our design leverages Hyperledger Iroha to enable lightweight decentralized consensus for ID verification, thereby mitigating risks like Sybil attacks in platooning. A malicious entity attempting to impersonate a platoon leader would be immediately detected, isolated, and denied participation, ensuring safety and integrity. Collectively, the design addresses CA centralization, revocation latency, and linkability risks identified in prior work [12, 13, 11, 7, 5, 14].
@@
Our framework is validated through both conceptual analysis and preliminary simulations. The simulations, spanning urban and highway scenarios, show minimal impact on network performance while ensuring robust identity management. Specifically, authentication introduces negligible delay and only minor overhead in communication metrics, demonstrating practicality for real-world deployment. Our evaluations in urban and highway scenarios using C-V2X technology indicate practical overheads: less than 7% reduction in Packet Reception Ratio (PRR) in urban settings, 29% in highway settings (300 vehicles/km), under 10% increase in Channel Busy Ratio (CBR), and under 0.06 s added latency.
@@
  • Zero Trust–based ID management. A blockchain-enabled, decentralized identity layer that continuously verifies entities and constrains privileges to the least necessary scope.
@@
  • Testing scenario. A realistic urban/highway simulation setup reflecting dense, dynamic vehicular conditions and cooperative applications (e.g., platooning).
  • Results and implications. Minimal overhead with strong authenticity and accountability guarantees: PRR degradation under 7% (urban) and 29% (highway, 300 veh/km), CBR increase under 11%, and sub–0.06 s added latency when C-V2X used.
@@
The remainder of this paper is structured as follows. Section 2 reviews related work. Section 3 details the proposed D-IM system architecture and protocol. Section 4 presents performance evaluations and comparisons. Section 5 concludes and outlines future directions.
@@
2 RELATED WORK
Research on securing autonomous vehicles (AVs) and connected transportation systems spans multiple domains, from connected vehicle applications and cybersecurity threats to identity management. A broad body of work has examined the benefits of V2X applications, the vulnerabilities of increased connectivity, and the limitations of existing PKI-based identity frameworks. More recently, Zero Trust Architectures (ZTA) have been proposed as a paradigm shift from perimeter-based defenses, while blockchain has been investigated for its potential to provide decentralized trust, auditability, and tamper-resistance in vehicular networks. In this section, we review related work across these themes, highlighting key advances and limitations.
2.1 CONNECTED VEHICLE APPLICATIONS
Connected vehicles are central to the intelligent transportation systems (ITS) evolution, enabling a broad range of applications extending beyond simple automation of driving tasks. Early visions of Vehicular Ad Hoc Networks (VANETs) anticipated the transition from traditional vehicular networking toward advanced cooperative systems supporting safety, efficiency, and emerging industrial applications [15].
The evolution of connected and autonomous vehicles enables cooperative applications that transform mobility and logistics. For instance, platooning allows multiple vehicles to travel in coordinated convoys with reduced inter-vehicle spacing, which improves fuel efficiency, road capacity, and safety. Early works demonstrated the feasibility of platoon management through VANET-enabled cooperative adaptive cruise control [16], while more recent efforts emphasize eco-driving strategies at signalized intersections to enhance sustainability [17]. Beyond platooning, connected vehicles enable cooperative perception, where vehicles share sensor data to extend their situational awareness beyond the line-of-sight of individual sensors. This supports advanced safety functions like collision avoidance and cooperative maneuvers, which are critical for applications such as emergency vehicle prioritization and logistics optimization [18, 19].
The realization of these applications relies on robust V2X communication protocols. Dedicated Short-Range Communications (DSRC) has traditionally been used in field trials, while 5G/Cellular V2X (C-V2X) is increasingly adopted for its higher bandwidth, lower latency, and broader ecosystem support. Recent studies have also investigated the coexistence of DSRC and C-V2X in the 5.895–5.925 GHz spectrum, providing performance evaluations of these technologies under shared spectrum conditions [20, 21]. Together, these protocols provide the foundation for cooperative perception, platooning, and other connected vehicle applications, forming the backbone of ITS
2.2 AV CYBERSECURITY THREATS AND CHALLENGES
Despite the transformative potential of connected and autonomous vehicles, they expose systems to a broad and evolving set of cybersecurity threats. Early demonstrations, such as the Jeep Cherokee and Tesla Model S exploits, highlighted how attackers could remotely manipulate safety-critical systems, including braking and steering, through vulnerabilities in infotainment and controller area network (CAN) bus interfaces [8]. These incidents demonstrated the insufficiency of traditional perimeter-based defenses and established cybersecurity as a foundational requirement for safe AV deployment [9, 5].
AVs face diverse attack surfaces spanning sensors, in-vehicle networks, wireless communication links, and machine learning subsystems. Early works exploited weaknesses in a vehicle’s Controller Area Network (CAN bus) [22] and demonstrated how attackers could compromise a 2014 Jeep Cherokee by reprogramming a gateway chip to access critical subsystems [23, 24]. Similar vulnerabilities have been found in other vehicles, such as the Toyota Prius and Ford Escape, by targeting Electronic Control Units (ECUs) and head units [25, 8]. Beyond the CAN bus, AVs are susceptible to attacks on navigation systems, such as GPS receivers, which can lead to misrouting or vehicle hijacking [26, 27].
At the sensor layer, spoofing/jamming pose major threats. GPS signals can be spoofed/jammed, leading to misrouting or vehicle immobilization. LiDAR, radar, and camera are vulnerable to occlusion, reflection, or adversarial perturbations. Recent studies demonstrate physical-world spoofing and removal attacks on LiDAR-based perception systems [28, 29], while surveys further review sensor failures and spoofing/jamming risks across AV perception stacks [30]. At the in-vehicle network layer, the CAN bus remains widely deployed, but vulnerable to injection, replay, and denial-of-service (DOS) attacks due to its lack of built-in authentication [31, 32]. At the wireless communication layer, V2X links can be disrupted by intelligent jamming and DOS attacks [33, 34], which degrade cooperative applications like platooning and cooperative perception. The integration of machine learning into perception and intrusion detection introduces new adversarial risks, where carefully crafted perturbations can mislead models for traffic sign recognition, lane detection, or anomaly detection [35, 36, 37].
Survey efforts provide valuable taxonomies of these threats. Hussain and Zeadally present an early comprehensive survey of AV security challenges [38], while Amoozadeh et al. highlight vulnerabilities in cooperative driving streams [39], and Luo and Hou discuss countermeasures for intelligent and connected vehicles [40]. More recent surveys offer updated perspectives, including Sedar et al. on V2X cybersecurity mechanisms [41], Kifor et al. on automotive cybersecurity frameworks and monitoring [42], and Durlik et al. on readiness for AV cybersecurity [43]. Collectively, these works confirm that AVs face a multi-layered threat landscape that continues to evolve, underscoring the need for multi-layered defenses that integrate anomaly detection, resilient communication protocols, and secure, continuous, and decentralized security frameworks. Identity management emerges as a cornerstone for securing V2X communications, ensuring that only authenticated and verified entities can participate in safety-critical data exchange.
2.3 IDENTITY MANAGEMENT IN V2X COMMUNICATIONS
Identity management is a cornerstone of secure V2X communications, ensuring that only authenticated entities can participate in the exchange of safety-critical messages. Formal frameworks have been established by standardization bodies like IEEE Std 1609.2 [44] and ETSI TS 103 097 [45], which form the basis of Public Key Infrastructure (PKI)-based vehicular credential management systems. In this approach, vehicles obtain long-term certificates from a root authority and use short-term pseudonym certificates to preserve privacy and unlinkability [46]. However, this anonymity comes at a cost, as regulatory and forensic requirements often necessitate conditional traceability, creating a tradeoff between privacy and accountability [47, 46, 48].
A central challenge in PKI-based V2X is certificate revocation. Traditional methods like Certificate Revocation Lists (CRLs) require frequent, bandwidth-intensive updates, which are impractical in highly mobile vehicular networks. Lightweight schemes have been proposed to mitigate this issue [49], but ensuring timely distribution of revocation information remains difficult in dense V2X environments [50, 51]. Scalability and privacy are also ongoing limitations. High pseudonym change rates strain storage and distribution systems, while insufficient changes increase linkability risks [48]. Recent surveys emphasize that current PKI-based identity frameworks struggle to balance these needs [6, 48, 52, 53, 6, 42]. These limitations highlight the need for more robust, decentralized approaches to identity management, motivating the exploration of blockchain and Zero Trust principles as complementary paradigms.

### v2

@@ Abstract
Autonomous vehicles (AVs) rely on pervasive connectivity to enable cooperative and safety-critical applications, but this connectivity also exposes them to a wide range of cybersecurity threats. Existing perimeter-based security and centralized identity management approaches are inadequate for highly dynamic V2X environments, as they depend on implicit trust and suffer from scalability and single-point-of-failure limitations. This paper proposes D-IM, a Zero Trust-based decentralized identity management and authentication framework for secure V2X communication. D-IM integrates continuous verification with a permissioned blockchain to eliminate centralized trust assumptions and enforce explicit, verifiable identity relationships among vehicles and infrastructure. The framework is designed around clear Zero Trust-aligned goals, including mutual authentication, decentralization, privacy protection, non-repudiation, and traceability, and addresses a comprehensive attacker model covering identity, data integrity, collusion, availability, and accountability threats. We present the D-IM system architecture and identification and authorization protocol, and validate its security properties through both qualitative analysis and a formal BAN logic-based verification. Simulation results in urban and highway scenarios using DSRC and C-V2X demonstrate that D-IM introduces limited overhead while preserving network performance, supporting its practicality for real-world AV deployments.
keywords:
@@ 1 INTRODUCTION
Autonomous Vehicles (AVs) are revolutionizing modern transportation by offering improved road safety, efficiency, and convenience, while minimizing human error that is responsible for more than 90% of road accidents and more than 40,000 fatalities and 2 million injuries annually in the United States alone [singh2015critical, national2020early]. AVs integrate sophisticated sensors, Electronic Control Units (ECUs), and advanced communication protocols to enable semi- and fully autonomous driving. These systems rely on Global Positioning Systems (GPS), LiDAR, cameras, and Vehicle-to-Everything (V2X) communication for perception, situational awareness, and decision-making [bhatia2019autonomous].
The Society of Automotive Engineers (SAE) defines six levels of driving automation, ranging from Level 0 (no automation) to Level 5 (full automation). These levels describe the degree of vehicle autonomy in terms of steering control, environment monitoring, and fallback responsibility [satam2022autonomous, yousseef2025autonomous]. Table 1 summarizes the SAE autonomy levels. As AVs progress toward higher autonomy (aiming for Level 5), their reliance on connectivity becomes indispensable. V2X communication, including Vehicle-to-Vehicle (V2V), Vehicle-to-Infrastructure (V2I), and Vehicle-to-Network (V2N) enables cooperative mobility. Key use cases include platooning, where vehicles form convoys to reduce drag and improve fuel efficiency; drone swarms, coordinating with AVs for delivery or traffic monitoring; emergency vehicle prioritization, where AVs yield dynamically to ambulances and police; and cooperative perception, where vehicles share sensor data for beyond-line-of-sight awareness [ying2024literature]. Dedicated Short-Range Communications (DSRC) and cellular-V2X (C-V2X) are the primary enabling radio access technologies, supporting low-latency and high-reliability communication.
@@ 5 Full S S S All
In platooning, for instance, the lead vehicle periodically broadcasts kinematic states (e.g., position, speed, acceleration) and intent, while the follower vehicles fuse these messages with onboard sensors to compute control inputs that maintain inter-vehicle gaps and string stability where the disturbances are not amplified when propagating along the vehicles string. While this cooperative driving strategy improves safety and fuel efficiency, it also introduces significant vulnerabilities. The increased interconnection introduces severe cybersecurity concerns. For instance, malicious actors can impersonate a leader vehicle, injecting false acceleration or braking commands that endanger the convoy. Once inside, the adversary could launch attacks such as fake data injection, transmitting falsified distance or braking information that destabilizes the convoy, or GPS spoofing, misleading the platoon about its true location and causing unsafe maneuvers. These manipulations not only threaten the cohesion of the platoon but also heighten the risk of chain-reaction accidents. Such scenarios highlight the necessity of robust identity verification and trust management mechanisms to safeguard cooperative vehicular systems.
The increased connectivity and complexity of AVs increases AVs cybersecurity vulnerabilities introducing many attack vectors. Attack vectors such as GPS spoofing, sensor manipulation, V2X spoofing and information injection threaten vehicle safety and operational integrity. Real-world incidents, such as the Jeep Cherokee and Tesla Model S exploits, have demonstrated the potential for remote hijacking and malicious manipulation of vehicle systems [miller2015remote]. These incidents highlight the inadequacy of traditional perimeter-based security models, which assume trust within a protected network, and the urgent need for robust multilayered defense strategies [anderson2023zero]. Unfortunately, many traditional approaches rely on trusted entities for credential management, a critical weakness adversaries can exploit by stealing or maliciously issuing credentials.
Identity management thus becomes critical for validating trust in dynamic vehicular environments. Current solutions primarily rely on Public Key Infrastructure (PKI) and pseudonym certificates, but these approaches face significant challenges in credential revocation, often failing to achieve timely and efficient invalidation of compromised credentials, thereby exposing vehicles to prolonged risk [ndss2024Revocation]. Recent research highlights that secure and scalable identity management remains one of the most pressing V2X research challenges [ying2024literature, yousseef2025autonomous].
To secure AVs against the previously mentioned attacks, Zero Trust Architecture (ZTA) addresses limitations of perimeter models by eliminating implicit trust and enforcing continuous verification of all entities, whether inside or outside the network. ZTA aligns with AV needs where real-time decisions depend on authenticated, accurate data, and is advocated by NIST for modern, interconnected systems [nist800207]. Within this paradigm, identity management emerges as the core security enabler for V2X. Ensuring that only authenticated, authorized, and non-revoked entities can publish or act upon safety messages is essential, especially in dense, dynamic traffic. Existing Identity management approaches suffer from centralization, scalability bottlenecks, and slow or bandwidth-heavy revocation mechanisms where a single trusted authority with implicit trust is needed presention a potential single point of failure [ying2024v2xReview, ndss2024Revocation].
@@
This paper presents a Zero Trust-based Decentralized Identity Management (D-IM) system for AVs, instantiated on a permissioned blockchain. Our solution eliminates dependence on centralized authorities, supports (through the use of blockchain technology) low-latency identity proofing that meets V2X timing budgets, enables fast and tamper-evident revocation propagation across dynamic topologies, and scales to dense traffic scenarios while preserving privacy through unlinkable pseudonym use. Moreover, D-IM leverages blockchain-backed identifiers and verifiable credentials to bind cryptographic identities to vehicles and infrastructure, enforces continuous verification at enrollment and message time, and supports fast, tamper-evident revocation. [yousseef2025autonomous, nist800207] Conceptually, every entity, vehicle, sensor, or infrastructure node, must prove authenticity before joining the network. Identity proofs accompany safety messages so consumers can validate origin, freshness, and authorization before actuation. The decentralized ledger provides immutable auditability and eliminates implicit trust in perimeter boundaries, while Zero Trust policies restrict privileges to the minimum necessary and re-evaluate them continuously. Practically, our design leverages Hyperledger Iroha to enable lightweight decentralized consensus for ID verification, thereby mitigating risks like Sybil attacks in platooning. A malicious entity attempting to impersonate a platoon leader would be immediately detected, isolated, and denied participation, ensuring safety and integrity. Collectively, the design addresses CA centralization, revocation latency, and linkability risks identified in prior work [he2022survey, miglani2023blockchain, ying2024v2xReview, ndss2024Revocation, yousseef2025autonomous, vangala2020blockchain].
@@
We conduct both a qualitative and a formal security analyses using Burrows-Abadi-Needham (BAN) logic to demonstrate that the proposed framework effectively addresses the attacks defined in the threat model. We validate our framework through both conceptual analysis and preliminary simulations. The simulations cover both urban and highway scenarios for two separate radio access technologies, namely Dedicated Short Range Communications (DSRC) and Cellular Vehicle-to-Everything (C-V2X). These simulations show minimal impact on network performance while ensuring robust identity management. Our evaluations in urban and highway scenarios using C-V2X technology indicate practical overheads: less than 19% reduction in Packet Reception Ratio (PRR) in urban settings, 3% in highway settings (200 vehicles/km), under 16% increase in Channel Busy Ratio (CBR), and under 0.032 s added latency. Our framework’s authentication introduces negligible delay and only minor overhead in communication metrics, demonstrating practicality for real-world deployment
@@
  • We propose Zero Trust-based Blockchain-enabled ID Management and Authentication for AVs.
  • We present a qualitative and formal BAN logic-based security analysis demonstrating robustness against the considered threat model.
@@
  • We perform a realistic urban/highway simulations reflecting dense, dynamic vehicular conditions and cooperative applications
@@
The remainder of this paper is structured as follows. Section 2 reviews related work. Section 3 presents the D-IM design goals, system and attacker models. Section 4 details the proposed D-IM system architecture and protocol. Section 5 provides a formal and qualitative security analysis. Section 6 presents performance evaluations and comparisons. Section 7 concludes and outlines future directions.
2 RELATED WORK
Research on securing autonomous vehicles and connected transportation systems spans multiple domains, from connected vehicle applications and cybersecurity threats to identity management. A broad body of work has examined the vulnerabilities of increased connectivity, and the limitations of existing identity management frameworks. More recently, Zero Trust Architectures (ZTA) have been proposed as a paradigm shift from perimeter-based defenses, while blockchain has been investigated for its potential to provide decentralized trust, auditability, and tamper-resistance in vehicular networks. In this section, we review related work across these themes, highlighting key advances and limitations in: (1) AV cybersecurity threats and challenges, (2) Zero Trust principles and evolution, (3) Zero Trust and blockchain integration, and (4) blockchain in vehicular and ITS security.
2.1 AV Cybersecurity Threats and Challenges
Despite the transformative potential of connected and autonomous vehicles, they expose systems to a broad and evolving set of cybersecurity threats. Early demonstrations, such as the Jeep Cherokee and Tesla Model S exploits, highlighted how attackers could remotely manipulate safety-critical systems, including braking and steering, through vulnerabilities in infotainment and controller area network (CAN) bus interfaces [miller2015remote]. These incidents demonstrated the insufficiency of traditional perimeter-based defenses and established cybersecurity as a foundational requirement for safe AV deployment [anderson2023zero, yousseef2025autonomous].
AVs face diverse attack surfaces spanning sensors, in-vehicle networks, wireless communication links, and machine learning subsystems. Early works exploited weaknesses in a vehicle’s CAN bus [hoppe2008security] and demonstrated how attackers could compromise a 2014 Jeep Cherokee by reprogramming a gateway chip to access critical subsystems [miller2014survey, miller2019lessons]. Similar vulnerabilities have been found in other vehicles, such as the Toyota Prius and Ford Escape, by targeting Electronic Control Units (ECUs) and head units [checkoway2011comprehensive, miller2015remote]. Beyond the CAN bus, AVs are susceptible to attacks on navigation systems, such as GPS receivers, which can lead to misrouting or vehicle hijacking [psiaki2016attackers, abrar2024gps].
At the sensor layer, spoofing and jamming pose major threats. GPS signals can be spoofed/jammed, leading to misrouting or vehicle immobilization. LiDAR, radar, and camera sensors are vulnerable to occlusion, reflection, or adversarial perturbations. Recent studies demonstrate physical-world spoofing and removal attacks on LiDAR-based perception systems [cao2023you, sato2025realism], while surveys further review sensor failures and spoofing/jamming risks across AV perception stacks [matos2024survey]. At the in-vehicle network layer, the CAN bus remains widely deployed, but vulnerable to injection, replay, and denial-of-service (DOS) attacks due to its lack of built-in authentication [al2024can, lampe2024can]. At the wireless communication layer, V2X links can be disrupted by intelligent jamming and DOS attacks [twardokus2023toward, arif2024clustered], which degrade cooperative applications like platooning and cooperative perception. Furthermore, the integration of machine learning into perception and intrusion detection systems introduces new adversarial risks, where carefully crafted perturbations can mislead models for traffic sign recognition, lane detection, or anomaly detection [chahe2023dynamic, pavlitska2023adversarial, kim2024survey].
Survey efforts provide valuable taxonomies of these threats. Hussain and Zeadally present an early comprehensive survey of AV security challenges [hussain2018autonomous], while Amoozadeh et al. highlight vulnerabilities in cooperative driving streams [amoozadeh2015security], and Luo and Hou discuss countermeasures for intelligent and connected vehicles [luo2019cyberattacks]. More recent surveys offer updated perspectives, including Sedar et al. on V2X cybersecurity mechanisms [sedar2023comprehensive], Kifor et al. on automotive cybersecurity frameworks and monitoring [kifor2024automotive], and Durlik and Dziubecki on readiness for AV cybersecurity [durlik2024cybersecurity]. Collectively, these works confirm that AVs face a multi-layered threat landscape that continues to evolve, underscoring the need for multi-layered defenses that integrate anomaly detection, resilient communication protocols, and secure, continuous, and decentralized security frameworks. Identity management emerges as a cornerstone for securing V2X communications, ensuring that only authenticated and verified entities can participate in safety-critical data exchange.